# Hedgehog topological defects in 3D amorphous solids

### v1

@@ Abstract
In Fig. S2, we show a snapshot of a 3D eigenvector field presented two clear topological defects with opposite charge.
@@
In Fig. S3(a), we present the vibrational density of states (vDOS) of the polymer glass model at $T=0$. As highlighted in the inset, at low frequency the vDOS exhibits a distinct Debye scaling, $D(\omega)\propto\omega^{2}$. In panel (b) of the same figure, we present the number of hedgehog topological defects in the eigenvector field $N_{\text{HTD}}$ as a function of frequency $\omega$. At low frequency, the number of defects also scales as $\omega^{2}$, implying a correlation with the vDOS.
@@
Figure S3: (a) The vibrational density of states $D(\omega)$ of the polymer glass model at $T=0$ extracted from the direct diagonalization of the Hessian matrix. The inset highlights the Debye scaling of the vDOS at low frequency. (b) The number of hedgehog topological defects in the eigenvector field $N_{\text{HTD}}$ as a function of frequency $\omega$. The inset zooms on the low frequency region and highlights the $\omega^{2}$ scaling.
Calculation and interpolation of non-affine displacement field on a cubic lattice
We derive the displacement field as Baggioli et al. (2021); Bera et al. (2024) $\vec{u}_{i}={\vec{r_{i}}}(\gamma)-{\vec{r}_{i}}(\gamma-\delta\gamma)$, where $\vec{r}_{i}(\gamma)$ and $\vec{r}_{i}(\gamma-\delta\gamma)$ represents the locations of the particle $i$ at strains $\gamma$ and $\gamma-\delta\gamma$, respectively. The affine component for the simple shear scenario is calculated as $\vec{u}_{A}=z\delta\gamma\hat{x}$. Finally we obtain the non-affine displacement field by subtracting the affine component from the total displacement, $\vec{u}_{\text{NA}}=\vec{u}_{i}-\vec{u}_{A}$ Baggioli et al. (2021); Bera et al. (2024). To obtain HTDs within the non-affine displacement field, we interpolated the discrete non-affine displacement field $\vec{u}_{\text{NA}}$ onto a simple cubic lattice grid with a side length $\sigma/2$ using the standard Python package griddata with linear interpolation.
@@
Appendix A Extended data and correlation between soft spots identified with different methods
@@
In Fig. S4, we provide extended data for the spatial correlation between between soft spots (s) identified using the softness parameter, Eq. (6), and negative (n) and positive (p) HTDs. Low-frequency data are provided in Fig. 2 in the main text. From Fig. S4, the vanishing of the correlation at high frequency ($\omega\gtrapprox 2$) is evident.
@@
Finally, for completeness, in Fig. S5 we show the spatial correlation between the plastic soft spots identified with method I (softness parameter, Eq. (6)) and method II (non-affine displacement field). From there, we observe that the two definitions provide a strong correlation at short distance that, nevertheless, does not survive for large length scales.
@@
Figure S4: Extended data showing the spatial correlation between soft spots (s) identified using the softness parameter, Eq. (6), and negative (n) and positive (p) HTDs for several higher frequencies. See Fig. 2 in the main text for low-frequency data.
@@
Figure S5: Spatial correlation between soft spots identified using the softness parameter, Eq. (6), and the non-affine displacement field. Different lines correspond to different values of strain, located around major plastic events along the stress-strain curve, Fig. S1(c).

### v2

@@ Acknowledgements
In Figure 8a, we present the vibrational density of states (vDOS) of the polymer glass model at $T=0$. As highlighted in the inset, at low frequency the vDOS exhibits a distinct Debye scaling, $D(\omega)\propto\omega^{2}$. In panel b of the same figure, we present the number of hedgehog topological defects in the eigenvector field $N_{\text{HTD}}$ as a function of frequency $\omega$. At low frequency, the number of defects also scales as $\omega^{2}$, implying a correlation with the vDOS.
@@
Figure 8: The vDOS and the number of HTDs versus eigenfrequencies. a The vibrational density of states $D(\omega)$ of the polymer glass model at $T=0$ extracted from the direct diagonalization of the Hessian matrix. The inset highlights the Debye scaling of the vDOS at low frequency. b The number of hedgehog topological defects in the eigenvector field $N_{\text{HTD}}$ as a function of frequency $\omega$. The inset zooms on the low frequency region and highlights the $\omega^{2}$ scaling.
Figure 9: 3D defect structures for different $N_{s}$. a–f, Snapshots of the 3D vector field on cubic cells. Defects with topological charge $Q=+1$ ($Q=-1$) at the center are shown in red (blue). The winding numbers $q_{2D}$ of the projected vector field on the cube’s faces are marked with stars: $q_{2D}=+1$ in green, $q_{2D}=-1$ in orange, and $q_{2D}=0$ in black. Each face is colored according to the corresponding 2D point defect. The panels illustrate various configurations with different values of $N_{s}$, the number of faces exhibiting $q_{2D}=-1$.
Figure 10: Behavior of HTDs with different saddle-surfaces. a The number of hedgehog topological defects, $N_{\text{HTD}}(N_{s})$, as a function of eigenfrequency $\omega$, for different values of $N_{s}$ (the number of saddle-surfaces). Data are shown in distinct colors for $N_{s}=0$ (blue), $N_{s}=1$ (magenta), $N_{s}=2$ (maroon), $N_{s}=3$ (green) and $N_{s}=4$ (indigo). b The fraction of HTDs, $\zeta=N_{\text{HTD}}(N_{s})/N_{\text{HTD}}$ plotted against $\omega$, for radial ($N_{s}=0$) and hyperbolic ($N_{s}>0$) defects.
Hyperbolic and radial HTDs with varying saddle-surfaces
To classify hyperbolic and radial defects, we project the three-dimensional vector field onto multiple two-dimensional planes and compute the winding number, $q_{2D}$, on the six faces of a cubic cell. The winding number takes discrete values of $+1$, 00, and $-1$, corresponding to different topological structures. In Figure 9, we present several 3D defect structures along with their 2D projections on the faces of cubic cells. Each face of the cube is color-coded based on its $q_{2D}$ value: green for $q_{2D}=+1$, black for $q_{2D}=0$, and orange for $q_{2D}=-1$. We define $N_{s}$ as the number of cube faces having $q_{2D}=-1$, representing the number of saddle-like surfaces. Here, we show several vector field configurations on cubic cells, highlighting different values of $N_{s}$ and their corresponding defect structures.
Behavior of HTDs with different $N_{s}$ as a function of eigenfrequency
In Figure 10a, we analyze the frequency dependence of the number of HTDs, $N_{\text{HTD}}(N_{s})$, for different values of $N_{s}$. Our results indicate that $N_{\text{HTD}}(N_{s})$ decreases with increasing $N_{s}$, suggesting that defects with multiple saddle-like projections occur less frequently. We compute the ratio $\zeta$, which represents the fraction of HTDs with a given $N_{s}$ relative to the total number of HTDs. In Figure 10b, we present this ratio separately for hyperbolic defects ($N_{s}>0$) and radial defects ($N_{s}=0$). At low frequencies, the difference in the number of hyperbolic and radial defects is small, while at higher frequencies, hyperbolic defects become more prevalent than radial defects.
Correlation between soft Spots and HTDs with hyperbolic and radial nature
We analyze the average correlation between soft spots and hyperbolic (H) defects, as well as soft spots and radial (R) defects, for different frequency cutoffs ($\omega_{c}$), as shown in Figure 11a. Our results indicate that the correlation with hyperbolic defects is consistently stronger than with radial defects across all $\omega_{c}$. Furthermore, we observe a systematic decrease in the peak height with increasing $\omega_{c}$, suggesting that the correlations are strongest at lower frequencies. In Figure 11b, we present the average correlation between soft spots and positive (p) HTDs, as well as soft spots and negative (n) HTDs. No significant differences are observed between these cases, consistent with the discussion in the main text.
Figure 11: Average correlations between soft spots and HTDs. a Average spatial correlation between soft spots (s), identified using the softness parameter (Eq. (7)), and hyperbolic (H) and radial (R) defects for different frequency cutoffs: $\omega_{c}=2.0$, $2.5$, and $4.0$. The correlation with hyperbolic defects remains consistently stronger than with radial defects across all $\omega_{c}$. b, Average correlation between soft spots (s) and hedgehog topological defects (HTDs) with positive (p) and negative (n) topological charges, showing no significant difference between the two cases.
@@
Extended data and correlation between soft spots identified with different methods
@@
In Figure 12, we provide extended data for the spatial correlation between between soft spots (s) identified using the softness parameter, Eq. (7), and negative (n) and positive (p) HTDs. Low-frequency data are provided in Fig. 2 in the main text. From Figure 12, the vanishing of the correlation at high frequency ($\omega\gtrapprox 2$) is evident.
@@
Finally, for completeness, in Figure 13 we show the spatial correlation between the plastic soft spots identified with method I (softness parameter, Eq. (7)) and method II (non-affine displacement field). From there, we observe that the two definitions provide a strong correlation at short distance that, nevertheless, does not survive for large length scales.
@@
Figure 12: Extended data on soft spots and HTDs with positive and negative charge. The spatial correlation between soft spots (s) identified using the softness parameter, Eq. (7), and negative (n) and positive (p) HTDs for several higher frequencies. See Fig. 2 in the main text for low-frequency data.
@@
Figure 13: Correlations between soft spots using method I and method II. Spatial correlation between soft spots identified using the softness parameter, Eq. (7), and the non-affine displacement field. Different lines correspond to different values of strain, located around major plastic events along the stress-strain curve, Figure 6c.